texconfig(1) teTeX texconfig(1) NAME texconfig−configuresteTeX SYNOPSIS texconfig DESCRIPTION texconfigallowstoconfigureandmaintainteTeXinaneasyandconvenientmanner.Itusesthe dialogprogramtodrawniceinputboxes.Afteransweringthedialoges,theappropriatefilesare changedautomaticallyandallnecessarystepsareinvolvedtoapplythenewsettings. SEEALSO dialog(1),initex(1),mf(1),vi(1),EDITORenvironment-variable. AUTHOR ThomasEsser(te@informatik.uni−hannover.de) teTeX 09/94 1